OpenStudy (deercult):

What are the real and complex solutions of the polynomial equation? x^4-41x^2=-400

OpenStudy (mathstudent55):

Do a substitution. Let \(u=x^2\). Then you get \(u^2=x^4\). Now substitute \(u^2\) for \(x^4\) and \(u\) for \(x^2\) in the given equation. You will now have a quadratic equation in \(u\). Solve for \(u\). Can you try the above and show what you get?

OpenStudy (deercult):

\[u^{2}-41u=-400\] Hmm... \[u^{2}-41u+400=0\] Just making sure, am I off on the right foot?

OpenStudy (deercult):

Okay, here's what I have so far, using the quadratic formula and that equation. \[x=\frac{ (41)\pm \sqrt{(-41)^{2}-4(1)(200)} }{ 2(1) }\]

OpenStudy (deercult):

@mathstudent55

OpenStudy (mathstudent55):

You are very close. Just remember, this is not x. Also, c = 400, not 200. You are now solving the new equation for u. Can you get values for u from your expression?

OpenStudy (mathstudent55):

\(u^{2}-41u+400=0\) \(u = \dfrac{-(-41) \pm \sqrt{(-41)^2 - 4(1)(400)}}{2(1)} \)

OpenStudy (deercult):

Oops, this whole time I was doing it in the wrong place and got the numbers confused, sorry! Now I have: \[u=\frac{ 41\pm \sqrt{1681-1600}}{ 2 }\] 1681-1600 is 81, and the square root of 81 is 9, leaving, \[u=\frac{ 41\pm9 }{ 2 }\] so far, yes?

OpenStudy (mathstudent55):

Excellent. Just the same I got. \(u = \dfrac{41 \pm \sqrt{1681 - 1600}}{2} \) \(u = \dfrac{41 \pm \sqrt{81 }}{2} \) \(u = \dfrac{41 \pm 9}{2} \)

OpenStudy (mathstudent55):

Now what two numbers does that simplify to?

OpenStudy (deercult):

41-9 = 32 / 2 = 16 41+9 = 50 / 2 = 25 Making 16 and 25 the solutions, I think

OpenStudy (mathstudent55):

Correct. I have to go. I'll be back in 10 minutes. Remember that those numbers are what u is equal to. Now substitute back. Let x^2 equal each of those two numbers and solve the two equations for x. \(x^2 = 25\) or \(x^2 = 16\) Each equation will give you two answers. The solution to the original equation is all 4 answers.

OpenStudy (deercult):

\[x^{4}-41(25)+400=0\] \[x^{4}-625=0\] \[x^{4}=625\] \[x=\pm5\] Two answers down!

OpenStudy (mathstudent55):

No. Once you have \(x^2 = 25\), you simply do this \(x = \pm \sqrt{25}\) \(x = \pm 5\) Now do the same for \(x^2 = 16\) to get the other two answers.

OpenStudy (deercult):

OH. Well, at the very least, I did get it right. :'> I got +/- 4i for the other two!

OpenStudy (mathstudent55):

Why i? 16 is a positive number just like 25. \(x^2 = 25\) \(x = \pm 5\) \(x^2 = 16\) \(x = \pm 4\)

OpenStudy (deercult):

Oh, that one I did goof up a little on, then, oops. I might've made a symbol mistake in my calculator. ^^; You're an amazing helper, thank you for taking all this time to help me out!!

OpenStudy (mathstudent55):

Since we have been doing parts and discussing them, it's easy to lose track of the flow of the steps of a complete solution, so here is a complete solution as you'd do in homework or in a quiz. Here is the entire solution from beginning to end. \(x^4-41x^2=-400\) \(x^4 - 41x^2 + 400 = 0\) Let \(u = x^2\), then \(u^2 = x^4\): \(u^2 - 41u + 400 = 0\) \(u = \dfrac{-(-41) \pm \sqrt{(-41)^2 - 4(1)(400)}}{2(1)} \) \(u = 25 \) or \(u = 16\) Substitute x back: \(x^2 = 25\) or \(x^2 = 16\) \(x = \pm \sqrt{25} \) or \(x = \pm \sqrt{16} \) \(x = 5\) or \(x = -5\) or \(x = 4 \) or \(x = -4\)

OpenStudy (mathstudent55):

You're welcome, and thanks for the kind words.

OpenStudy (mathstudent55):

gtg, bye

OpenStudy (deercult):

Alright, be safe and have a good night!

OpenStudy (mathstudent55):

ty, you too